OpenStudy (anonymous):

The population of a village is 5000. It is increasing at the rate of 10% each year. At the end of 3 years, 1155 people migrate from the village. The net increase in population at the end of 3 years is??

ganeshie8 (ganeshie8):

\[ net\ increase = 5000e^{.1*3} - 5000 - 1155\]

OpenStudy (anonymous):

Options are (A) 27% (B) 30% (C) 9% (D) 10%

ganeshie8 (ganeshie8):

ok i think here the given rate is fixed. so there is no point using continuous compounding. just use the compound growth rate formula and compute. i dont remember the exact formula sorry

OpenStudy (kropot72):

Population at the end of 3 years before migration = 5000(1.1)^3 = 6655 Population at the end of 3 years after migration = 6655 - 1155 = 5500 Net increase = 5500 - 5000 = 500 Percentage net increase is\[\frac{500}{5000}\times 100=?%\]

OpenStudy (anonymous):

Two pipes A and B can separately fill a cistern in 24 and 32 minutes respectively. They started to fill a cistern together, but B is turned off after few minutes and A fills the rest of the cistern. If total time taken is 18 minutes, after how many minutes is tap B turned off?